OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle whose legs are 5 and squareroot(2). 2squareroot(3) 3squareroot(3) squareroot(29)

OpenStudy (anonymous):

its multiple choice (:

OpenStudy (anonymous):

hypothenus are found by Sqrt[x^2+y^2]=h

OpenStudy (anonymous):

Sqrt[5^2+(sqrt[2])^2] sqrt[27]

OpenStudy (anonymous):

sqrt[27]=3sqrt[3]

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\] Where a and b are the length of the legs and c is the length of the hypotenuse. Solve for c.

OpenStudy (anonymous):

sqrt (5^2 + 2)

OpenStudy (anonymous):

\[\sqrt{5^2+\sqrt{2}^2}=\sqrt{25+2}=\sqrt{27}\]\]

OpenStudy (anonymous):

and since \[27=3^3\] you have \[\sqrt{3^3}=3\sqrt{3}\]

OpenStudy (anonymous):

so answer is second one